Epidemiology

In a hypothetical study investigating the interaction between factor A and factor B and the risk of disease X; Incidence of disease in the absence of factor A and factor B = 2 per 100,000 population Incidence of disease if factor A is present and factor B absent = 14 per 100,000 population Incidence of disease if factor B is present and factor A absent = 10 per 100,000 population

1. Using the additive model of interaction, what is the expected incidence in persons who are exposed to both factors A and B if those persons experienced the risk resulting from the independent contribution of both factors?
Answer #1

Additive interaction is originally described by Rothman, as a deviation from additivity of the absolute effects of two risk factors.

The interaction can be measured based on the ways that risks are calculated, the modelling.

Risks are considered to act in a multiplicative way when a ratio is used.

Risks are considered to act in an additive way when the difference is used.

The interaction arises when the incidence of disease in the presence of two or more risk factors differs from the incidence expected to result from each of their individual effects.

Implications of Interaction

Synergism:

increases disease risk beyond expected as the persons with one exposure are more susceptible to another

exposure

Antagonism:

decreases disease risk beyond expected as the persons with one exposure are less susceptible to

another

The expected incidence of A and B = Attributable risk of A alone + attributable risk of B alone + baseline

Or

= Incidence of A alone + incidence of B alone – baseline

Incidence of A alone = 14/100,000

Incidence of B alone=10/100,000

Baseline (without A & B) = 2/100,000

Additive interaction = A +B – Base line

= 14 + 10 – 2 = 22 is positive and super additive

  

If the interaction is > 0 it is said to be positive or “super-additive”

If the interaction is < 0 it is said to be negative or “sub-additive”